### v1

Abstract
Threshold-Computation-in-the-Head (TCitH) and VOLE-in-the-Head (VOLEitH), two recent developments of the MPC-in-the-Head (MPCitH) paradigm, have significantly improved the performance of digital signature schemes in this framework.
In this note, we embed the restricted decoding problem, denoted by $\mathcal{E}$-SDP, within these frameworks. We propose a structurally simple modeling that achieves competitive signature sizes. Specifically, by instantiating the restricted decoding problem with the same hardness assumption underlying CROSS, we reduce sizes by more than a factor of two compared to the NIST submission. Moreover, we observe that ternary full-weight decoding, closely related to the hardness assumption underlying WAVE, is a restricted decoding problem. Using ternary full-weight decoding, we obtain signature sizes comparable to the smallest MPCitH-based candidates in the NIST competition.
[…]
1 Introduction
Following the standardization of the signatures CRYSTALS-Dilithium, FALCON, and SPHINCS+, NIST issued a call for additional digital signature schemes in September 2022. This initiative, to which 40 proposals were submitted, has the intention of diversifying the portfolio of post-quantum signature standards. In October 2024, NIST announced that fourteen submissions are going to proceed to round two. Among these, six are based on the Threshold Computation in-the-Head (TCitH) and the VOLE in-the-head (VOLEitH) frameworks [1, 2, 3, 5, 12, 14]. These recent specializations of the Multi-Party Computation (MPC) in-the-Head (MPCitH) paradigm offer significant performance improvements compared to previous constructions.
1.0.1 MPCitH.
Originally proposed in 2007 by Ishai, Kushilevitz, Ostrovsky, and Sahai [23], the MPCitH techniques have reshaped the landscape of post-quantum digital signatures. In a nutshell, the MPCitH approach constructs digital signatures by combining secure multi-party computation (MPC) with the Fiat–Shamir transform. Emulating an MPC protocol, the prover (i.e., the signer) performs computations on shares of a secret. A zero-knowledge proof of knowledge is obtained by opening a verifier-selected subset of the parties, which is checked for consistency.
1.0.2 TCitH and VOLEitH.
The TCitH framework [21] instantiates the MPCitH paradigm with an $\ell$-out-of-$n$ threshold secret sharing. Alternatively, the TCitH framework may be described in the Polynomial Interactive Oracle Proofs (PIOP) formalism, as presented in [20]. This formalism also encompasses the VOLEitH framework, originally introduced as a technique to compile VOLE-based zero-knowledge protocols into publicly verifiable protocols, from which signatures can be obtained via the Fiat–Shamir transform [13]. Using the TCitH and VOLEitH frameworks, performant signatures can be obtained from various hardness assumptions, in particular from coding theory. Submissions to the NIST competition utilize Hamming-metric [3] and rank-metric decoding problems [2, 5], as well as the permuted kernel problem [1].
1.0.3 Our Contribution.
In this work, we consider the restricted decoding problem, called $\mathcal{E}$-SDP, underlying the CROSS signature scheme. We present a modeling of $\mathcal{E}$-SDP that allows constructing signatures in the TCitH and VOLEitH frameworks. While structurally simple, the proposed modeling achieves competitive performance. In fact, for the exact same decoding problem underlying CROSS, we reduce signature size by more than a factor of two. This shows that, by replacing the simple CVE-based proof of knowledge of CROSS with recent variants of the MPCitH framework, signature sizes can be reduced, albeit at the cost of a significantly more complex construction. Further, we consider the problem of ternary full-weight decoding, a special case of the hardness assumption underlying the WAVE signature scheme [10, 19]. Here, the proposed modeling achieves signature sizes that are comparable to the smallest ones of the NIST competition.
1.0.4 Organization.
Section 2.0.4 provides a concise review of the TCitH and VOLEitH frameworks. Table 3 presents the proposed modeling of $\mathcal{E}$-SDP, and discusses the achievable signature sizes.
1.0.5 Artifacts.
The accompanying repository [16] contains the code to reproduce the provided parameters.
2 Signatures from Threshold Computation in the Head
The VOLE-in-the-Head (VOLEitH) framework [13] and the Threshold-Computation-in-the-Head (TCitH) framework [21] are recent specializations of the MPC-in-the-Head paradigm [23]. Since we build our contribution in these frameworks, we provide a concise overview following the Polynomial Interactive Oracle Proof (PIOP) formalism presented in [20].
2.0.1 Notation.
For an integer $m\in\mathbb{Z}$, we denote $m]=\{1,\ldots,m\}$. Let $\mathbb{F}$ be the finite field with $p$ elements, where $p$ is a prime number, and $\mathbb{K}$ its degree-$\mu$ field extension. We denote by $\mathbb{F}[x_{1},\ldots,x_{n}]_{\leq d}$ the set of polynomials in variables $\mathbf{x}=(x_{1},\ldots,x_{n})$ with coefficients in $\mathbb{F}$ and total degree at most $d$. Vectors are denoted with bold lowercase letters, and matrices with bold uppercase letters. Polynomials are denoted in plain lowercase letters, and vectors of polynomials in plain uppercase letters. For a vector of polynomials $P(X)\in\mathbb{F}[X]_{\leq d}^{n}$, we denote by $P(\infty)\in\mathbb{F}^{n}$ the vector of leading (degree-$d$) coefficients.
2.0.2 TCitH in a Nutshell.
The general zero-knowledge proof of knowledge obtained from the TCitH framework is provided in Figure 1.
[…]
Figure 1: Five-pass zero-knowledge proof of knowledge.
The prover attempts to convince the verifier that they know a witness $\mathbf{w}\in\mathbb{F}^{n}$ satisfying $F\left(\mathbf{w}\right)=\mathbf{0}\in\mathbb{F}^{m}$ for a One-Way Function (OWF) $F$, described by the system of polynomial relations $F=(f_{1},\ldots,f_{m})\in\big(\mathbb{F}\left[x_{1},\ldots,x_{n}\right]_{\leq d}\big)^{m}$. To this end, the prover uses a GGM-based Vector Commitment (VC) scheme [11, 13] to sample vectors of random polynomials
$P_{u}\in\big(\mathbb{K}\left[X\right]_{\leq d-1}\big)^{\eta}\quad\text{and}\quad P_{w}\in\big(\mathbb{K}\left[X\right]_{\leq 1}\big)^{n}\quad\text{with}\quad P_{w}\left(\infty\right)\in\mathbb{F}^{n},$
which also generates a commitment $\mathsf{com}$ to $P_{w},P_{u}$ and the opening key $\mathsf{key}$. To embed the witness into $P_{w}$, the prover computes the offset $\boldsymbol{\Delta}_{\mathbf{w}}=\mathbf{w}-P_{w}\left(\infty\right)\in\mathbb{F}^{n}$, for which $\tilde{P}_{w}\left(X\right)=P_{w}\left(X\right)+\boldsymbol{\Delta}_{\mathbf{w}}\cdot X$ has leading [^1] Embedding the witness into the constant coefficient is also possible; see [14]. coefficients $\mathbf{w}$. The offset $\boldsymbol{\Delta}_{\mathbf{w}}$ and commitment $\mathsf{com}$ are then sent to the verifier. Next, the polynomial relation $F$ is evaluated at $\tilde{P}_{w}\left(X\right)$. Specifically, for $i\in[m]$ and $\ell\leq d$, denote by $f_{i,\ell}(\mathbf{x})$ the degree-$\ell$ homogeneous component of $f_{i}(\mathbf{x})$. Then, $F(\tilde{P}_{w})(X)$ is defined as
$F(\tilde{P}_{w})(X)\coloneqq\left(\sum_{\ell}X^{d-\ell}f_{1,\ell}\left(\tilde{P}_{w}\left(X\right)\right),\ldots,\sum_{\ell}X^{d-\ell}f_{m,\ell}\left(\tilde{P}_{w}\left(X\right)\right)\right).$
Then $F(\tilde{P}_{w})(\infty)=F\left(\mathbf{w}\right)$, i.e., the embedded witness satisfies the polynomial relation $F$ if and only if $F(\tilde{P}_{w})(X)$ has degree at most d1d-1. Hence, $F(\tilde{P}_{w})(X)$ consists of $m$ polynomials of degree d1d-1. The number of polynomials is reduced through batching. To this end, the verifier samples a random matrix $\boldsymbol{\Gamma}\in\mathbb{K}^{\eta\times m}$ and the prover computes $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})$, thereby reducing the number of polynomials from $m$ to $\eta$ while preserving the degree with overwhelming probability. The verifier masks $\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)$ by computing
$Q(X)=\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w})(X)+P_{u}\left(X\right),$
where $P_{u}\left(X\right)$ is the length-$\eta$ vector of random polynomials of degree (d1)(d-1) generated by the VC scheme. Sending the masked polynomial vector $Q(X)$ to the verifier does not reveal any information about the witness. To verify correctness, the verifier challenges the prover at a random evaluation point $r$ in the evaluation domain $\Omega\subseteq\mathbb{K}$ of size $N=\left\lvert\Omega\right\rvert$. The prover produces the corresponding opening $\mathsf{opening}_{r}$ using its opening key $\mathsf{key}$. The verifier checks $\mathsf{opening}_{r}$ against $\mathsf{com}$ and, in case of success, computes the evaluations $P_{w}\left(r\right)\in\mathbb{K}^{\eta}$ and $P_{u}\left(r\right)\in\mathbb{K}^{n}$. With these evaluations, the verifier computes $P_{u}\left(r\right)+\boldsymbol{\Gamma}\cdot F(\tilde{P}_{w}\left(r\right))$ and checks whether it matches $Q\left(r\right)$, where $Q(X)$ is the polynomial vector received from the prover.
Remark 1
In [14], also a three-pass variant of Figure 1 is given, which is obtained by skipping the batching step. This simplifies the protocol, but increases the required communication, which is why we omit this variant.
2.0.3 TCitH signature scheme.
The proof system above has a soundness error given by
$\varepsilon=\frac{1}{p^{\mu\cdot\eta}}+\left(1-\frac{1}{p^{\mu\cdot\eta}}\right)\frac{d}{N}.$
For a target security of $\lambda$ bits, $\tau$ parallel repetitions are performed to amplify soundness below $2^{-\lambda}$. The Fiat–Shamir transform (enhanced by a $w$-bit proof of work, known as grinding) yields a signature scheme with the following properties.
Property 1
For grinding parameter $w$, let the proof system parameters satisfy
$N\leq p^{\mu},\qquad p^{\mu\cdot\eta}\geq 2^{\lambda},\quad\text{and}\quad\left(\frac{N}{d}\right)^{\tau}\geq 2^{\lambda-w}.$
Then the resulting signature scheme is EUF-CMA-secure [14]. Signatures are of the form $\sigma=(\sigma_{\mathrm{sym}},\sigma_{\mathbf{w}},\sigma_{F})$, where $\sigma_{\mathbf{w}}$ corresponds to the witness, $\sigma_{F}$ to the OWF $F$, and $\sigma_{\mathrm{sym}}$ to the VC scheme. The components’ sizes [^2] In general, the sizes depend on the specific instantiation of the VC scheme [22], and optimizations not described here can be applied. The formulas given follow [14], to which we refer the reader for further details. are given by
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot\lambda\cdot(\log_{2}N+1)+5\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot\eta\cdot(d-1)\cdot\mu\cdot\log_{2}p.$
2.0.4 VOLEitH.
The VOLEitH framework is closely related to the TCitH framework described above. Instead of performing $\tau$ parallel repetitions, VOLEitH employs a technique that allows sampling $r$ from an evaluation domain $\Omega$ of size
$\left\lvert\Omega\right\rvert=N^{\tau}\approx 2^{\lambda}\leq\left\lvert\mathbb{K}\right\rvert=p^{\rho}.$
Avoiding parallel repetitions reduces the soundness error, which in turn enables shorter signatures. However, this comes at the cost of a consistency check, an additional round of interaction between the prover and the verifier. For further details, we refer to [12, 13].
By parameterizing the consistency check with $B$ and the state-of-the-art VC scheme [11] with $T_{\mathrm{open}}$, a signature scheme with the following properties can be constructed.
Property 2
Let the VOLEitH proof system be parameterized such that
$\frac{N^{\tau}}{d}\geq 2^{\lambda-w}\quad\text{and}\quad p^{\rho}\geq 2^{\lambda},\quad\text{with}\quad\rho=\tau\cdot\left\lceil\log_{p}N\right\rceil,\quad B=\left\lceil\frac{16}{\log_{2}p}\right\rceil.$
Then the resulting signature scheme is EUF-CMA-secure [12], and its signatures have size [^3] The provided formulas follow the specific design choices made in [3]. $\left\lvert\sigma\right\rvert=\left\lvert\sigma_{\mathrm{sym}}\right\rvert+\left\lvert\sigma_{\mathbf{w}}\right\rvert+\left\lvert\sigma_{F}\right\rvert$, where
$\left\lvert\sigma_{\mathrm{sym}}\right\rvert$ $=\tau\cdot(\rho+B)\log_{2}p+\tau\cdot 2\lambda+T_{\mathrm{open}}\cdot\lambda+4\lambda+32,$
$\left\lvert\sigma_{\mathbf{w}}\right\rvert$ $=\tau\cdot n\cdot\log_{2}p,$
$\left\lvert\sigma_{F}\right\rvert$ $=\tau\cdot(d-1)\cdot\rho\cdot\log_{2}p.$
Similar to the TCitH framework, which requires $\eta\cdot\mu\cdot\log_{2}p\geq\lambda$, the VOLEitH framework requires $\rho\cdot\log_{2}p\geq\lambda$. However, VOLEitH allows for a smaller $\tau$, owing to its reduced soundness error, arguably at the cost of a more complex scheme.
3 Modeling Restricted Decoding Problems
In this section, we present a so-called modeling for the restricted decoding problem, i.e., we express the OWF as a polynomial relation that can be used in the TCitH and VOLEitH frameworks. Before doing so, we briefly recap some basic notions from coding theory.
3.0.1 Coding theory.
A linear code over $\mathbb{F}$ of length $n$ is a $k$-dimensional subspace of $\mathbb{F}^{n}$. The code can be represented by a parity-check matrix $\mathbf{H}\in\mathbb{F}^{(n-k)\times n}$ such that $\mathbf{c}\mathbf{H}^{\top}=\mathbf{0}$ for every codeword $\mathbf{c}$. We denote by r=nkr=n-k the redundancy of the code. The syndrome of a vector $\mathbf{v}$ (through $\mathbf{H}$) is $\mathbf{s}=\mathbf{v}\mathbf{H}^{\top}\in\mathbb{F}^{r}$. The Hamming weight (from now on, only weight) of a vector $\mathbf{v}$ is the number of non-zero entries in $\mathbf{v}$.
3.0.2 Restricted decoding assumption.
The (general) restricted Syndrome Decoding Problem ($\mathcal{E}$-SDP) [8] replaces the Hamming-weight constraint of the classical SDP with the requirement that each entry must lie in the fixed subset $\mathcal{E}$ of the underlying finite field $\mathbb{F}$. Formally, it is defined as follows:
Definition 1($\mathcal{E}$-SDP)
Let $\mathcal{E}\subset\mathbb{F}$ denote the restriction, $\mathbf{H}\in\mathbb{F}^{r\times n}$ the parity-check matrix, and $\mathbf{s}\in\mathbb{F}^{r}$ the syndrome. The $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ is solved by $\mathbf{e}\in\mathcal{E}^{n}$ with $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$.
We refer to the size of the restriction as $z=\left\lvert\mathcal{E}\right\rvert$. Different choices of $\mathcal{E}$ are possible, and all of them lead to an NP-complete problem [9]. Restrictions of the form $\mathcal{E}=\{-a,\ldots,a\}$ for $a\ll p$ yield a problem similar in nature to learning with errors (LWE) [8], while [9] considers cyclic subgroups of $\mathbb{F}^{*}=\mathbb{F}\setminus\{0\}$. In general, any choice of restriction requires a dedicated analysis of the problem’s concrete computational hardness. Therefore, we focus on specific choices that have undergone careful analysis.
CROSS restriction.
The signature scheme CROSS [6] uses the restriction $\mathcal{E}=\{2^{i}\mid i\in[7]\}$ within $\mathbb{F}$ of size $p=127$. We are going to refer to this setting as CROSS-SDP, the concrete hardness of which is analyzed in detail in [7] and confirmed in [15]. For instance, the parameters $n=127$ and $k=76$ provide a suitable choice for NIST security category 1.
Ternary decoding with full weight.
The signature scheme WAVE [19, 10] relies on the ternary syndrome decoding with high weight. For state-of-the-art solvers [17], parameterizing the problem such that the error is full-weight and unique, maximizes the concrete hardness (a setting in which WAVE cannot sign successfully). This corresponds to $\mathcal{E}$-SDP with $\mathcal{E}=\{1,2\}$ and $p=3$, to which we are going to refer as Ternary-SDP. For $k\approx 0.369\cdot n$, the solver presented in [17] requires $\mathcal{O}\left(2^{0.247\cdot n}\right)$ operations, up to a polynomial factor. In particular, the parameters $n=579$ and $k=213$ are shown to achieve NIST security category 1.
3.0.3 Modeling $\mathcal{E}$-SDP.
The linearity of the syndrome equations allows a standard optimization that reduces the witness size from $n$ to $k$. We assume that the parity-check matrix is in systematic form [^4] This assumption, also made e.g. in [4], does not impact the hardness of $\mathcal{E}$-SDP. $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$, where $\mathbf{A}\in\mathbb{F}^{r\times k}$ and $\mathbf{1}_{r}\in\mathbb{F}^{r\times r}$ denotes the identity matrix. Then, $\mathbf{e}\in\mathbb{F}^{n}$ satisfies $\mathbf{e}\mathbf{H}^{\top}=\mathbf{s}$ if and only if it is of the form $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ with $\mathbf{w}\in\mathbb{F}^{k}$. That is, the witness is defined as a length-$k$ partial error vector that is expanded to a length-$n$ error vector so that the syndrome equations are satisfied.
[…]
Overall, we obtain the system of polynomial constraints $F(\mathbf{x})=(f_{1},\ldots,f_{n})\in\mathbb{F}[x_{1},\ldots,x_{k}]^{n}$ with
$f_{i}(\mathbf{x})=f_{\mathcal{E}}\left(x_{i}\right)\quad\text{for }i\in[k],\quad\text{and}\quad f_{k+i}(\mathbf{x})=f_{\mathcal{E}}\left(s_{i}-\langle\mathbf{a}_{i},\mathbf{x}\rangle\right)\quad\text{for }i\in[r],$ (1)
where $\mathbf{a}_{i}\in\mathbb{F}^{k}$ denotes the $i$-th row of $\mathbf{A}$. The correctness and degree of this modeling are stated in the following proposition.
Proposition 1
The polynomial relation $F$ given in Equation 1 provides a degree-$z$ modeling of $\mathcal{E}$-SDP: If $\mathbf{w}\in\mathbb{F}^{k}$ satisfies $F\left(\mathbf{w}\right)=\mathbf{0}$, then the error vector $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$ solves the $\mathcal{E}$-SDP instance $(\mathbf{H},\mathbf{s})$ with $\mathbf{H}=(\mathbf{A},\mathbf{1}_{r})$.
Proof
For $i\in[k]$, $f_{i}(\mathbf{x})$ has degree $z$. The composition of a degree-$z$ polynomial and a linear polynomial is also of degree $z$; and, thus, $f_{k+i}(\mathbf{x})$ and $F(\mathbf{x})$ have degree $z$ as well. Let $\mathbf{w}$ satisfy $F\left(\mathbf{w}\right)=\mathbf{0}$. Then, for $\mathbf{e}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})$, it holds
$\mathbf{e}\mathbf{H}^{\top}=(\mathbf{w},\mathbf{s}-\mathbf{w}\mathbf{A}^{\top})(\mathbf{A},\mathbf{1}_{r})^{\top}=\mathbf{w}\mathbf{A}^{\top}+\mathbf{s}-\mathbf{w}\mathbf{A}^{\top}=\mathbf{s},$
i.e., the syndrome equations are satisfied. Further, $f_{\mathcal{E}}\left(e_{i}\right)=0$ for all entries of $\mathbf{e}$, i.e., $\mathbf{e}\in\mathcal{E}^{n}$. ∎
3.0.4 Performance in TCitH framework.
From the provided modeling of $\mathcal{E}$-SDP, a digital signature is derived using the TCitH framework [21], as summarized in Section 2.0.4. The chosen parameters and the obtained signature sizes are provided in Table 1. The parameters $\tau$, $N$, and $w$ of the proof system are selected following [14] to allow for comparable runtime performance of the VC scheme. To simplify implementation, the smallest extension degree $\mu=[\mathbb{K}:\mathbb{F}]$ satisfying the requirement $N\leq\left\lvert\mathbb{K}\right\rvert=\left\lvert\mathbb{F}\right\rvert^{\mu}$ is chosen. This, in turn, determines the batching parameter $\eta$.
Table 1: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the TCitH framework. The proof system parameters are selected analogously to those in MQOM [14] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1
$127$
77
$127$
$76$
$24$ $256$ 22 $10$ 44 fast $7650$
$15$ $2048$ 22 $10$ 66 short $5533$
33
22
$579$
$213$
$17$ $256$ 66 $14$ 99 fast $3533$
$12$ $2048$ 77 $12$ 88 short $3095$
NIST 3
$127$
77
$187$
$111$
$36$ $256$ 22 $14$ 66 fast 16 67516\,675
$23$ $2048$ 22 $14$ 44 short 12 35412\,354
33
22
$839$
$309$
$27$ $256$ 66 2121 33 fast $8284$
$18$ $2048$ 77 $18$ $12$ short $6860$
NIST 5
$127$
77
$251$
$150$
$48$ $256$ 22 1919 77 fast 29 83929\,839
$31$ $2048$ 22 1919 33 short 22 30522\,305
33
22
$1102$
$406$
$36$ $256$ 66 $27$ 44 fast 14 58414\,584
$25$ $2048$ 77 $24$ 66 short 12 60812\,608
For Ternary-SDP, parameterized to achieve NIST security category 1, a signature size as small as $3.1\text{\,}\mathrm{kB}$ is possible. Switching to CROSS-SDP, the signature size increases to 5.5 $5\text{\,}\mathrm{kB}.$ This is mainly due to the increased degree of the modeling and a larger witness size. Nevertheless, we obtain a significant reduction compared to CROSS [6], for which the smallest signature size is $12.4\text{\,}\mathrm{kB}$.
3.0.5 Performance in VOLEitH framework.
Table 2 provides an overview of the performances that can be achieved if TCitH is replaced by the VOLEitH framework. Parameters of the proof system are chosen analogously to SDitH [3]. For Ternary-SDP, only a minor improvement in signature size is achievable. A similar effect is observed in [14]: for OWFs with small degree and small witness size, the structurally simpler TCitH framework remains competitive. On the other hand, for CROSS-SDP, which has a degree of $d=z=7$, VOLEitH enables a reduction in signature size of about $1\text{\,}\mathrm{kB}$ for NIST security category 1, and of up to $6\text{\,}\mathrm{kB}$ for NIST security category 5.
Table 2: Parameters of signatures based on $\mathcal{E}$-SDP obtained via the VOLEitH framework. The proof system parameters are selected analogously to those in SDitH [3] to ensure comparable runtime performance of the VC scheme.
Security
category
$\mathcal{E}$-SDP Proof System Performance
[…]
NIST 1
$127$
77
$127$
$76$
$16$ $256$ $32$ $101$ 33 fast $6432$
$11$ $2048$ $22$ $107$ $10$ short $4372$
33
22
$579$
$213$
$16$ $256$ $96$ $101$ 11 fast $3515$
$11$ $2048$ $81$ $107$ 88 short $2974$
NIST 3
$127$
77
$187$
$111$
$24$ $256$ $48$ $153$ 33 fast 14 35914\,359
$16$ $4096$ $32$ $157$ 33 short $9361$
33
22
$839$
$309$
$24$ $256$ $144$ $153$ 11 fast $7816$
$16$ $4096$ $128$ $157$ 11 short $6463$
NIST 5
$127$
77
$251$
$150$
$32$ $256$ $64$ $206$ 33 fast 25 57325\,573
2121 $4096$ $42$ $216$ 77 short 16 58916\,589
33
22
$1102$
$406$
$32$ $256$ $192$ $206$ 11 fast 13 85113\,851
2121 $4096$ $168$ $216$ 55 short 11 52111\,521
3.0.6 Comparison with NIST submissions.
A comparison of the achieved signature sizes with submissions to NIST’s call for additional digital signature schemes [24] is provided in Table 3. As can be seen, the TCitH and VOLEitH frameworks can be combined with a wide range of OWFs. The proposed modeling of $\mathcal{E}$-SDP is competitive compared to the state-of-the-art: for the Ternary-SDP, we obtain a signature size which is very close to those of the smallest proposals that utilize the TCitH or the VOLEitH framework with an alternative hardness assumption. The sizes obtained from CROSS-SDP are slightly larger than those of SDitH and PERK, which also use modelings of degree greater than two, but smaller than $z=7$. The code-based proposals CROSS and LESS target other trade-offs: CROSS uses a simple CVE-like protocol [18, 6], and LESS achieves a small signature size by using multiple public keys (resulting in a public-key size that is multiple times larger than those of the other schemes shown).
Table 3: Comparison of the obtained signature sizes with NIST submissions. The shown sizes correspond to the short optimization tradeoff; runtime may vary.
Scheme Hardness Assumption Design Principle $\left\lvert\sigma\right\rvert$ in $\mathrm{kB}$]
this work CROSS-SDP VOLEitH $4.4$
Ternary-SDP TCitH $3.1$
CROSS CROSS-$\mathcal{E}$-SDP CVE $12.4$
LESS Code Equivalence sigma protocol $1.3$
SDitH Syndrome Decoding VOLEitH $3.7$
RYDE Rank Decoding TCitH $3.1$
Mirath MinRank TCitH $3.0$
PERK Permuted Kernel VOLEitH $3.5$
MQOM Multivariate Quadratic TCitH $2.8$
FEAST AES VOLEitH $4.5$
AES-EM $3.9$
4 Conclusion
This paper presents a polynomial modeling of the restricted decoding problem $\mathcal{E}$-SDP, enabling the construction of digital signatures in the TCitH and the VOLEitH frameworks. When instantiated with ternary full-weight decoding, we obtain signature sizes that are competitive with the state-of-the-art NIST candidates. For the restriction used by the CROSS signature scheme, the sizes are significantly reduced compared to CROSS.